The sum of three positive numbers is 90 and the sum of them is the third. Find the numbers if the smallest one is half of the largest.
I have found 2 of the equations. I just need help trying to find the third one
1. x + y + z = 90
2. x + y = z
3. ????

Found 2 solutions by macston, rothauserc:
Answer by macston(5194)

The sum of three positive numbers is 90 and the sum of two of them is the third. Find the numbers if the smallest one is half of the largest.
x+y+z=90
2x=z
x+y=z ; y=z-x ; y=2x-x=x
x+y+z=90
x+x+2x=90
4x=90
x=22.5 ANSWER 1: x=22.5
z=2x=45 ANSWER 2: z=45
y=x=22.5 ANSWER 3: y=22.5

Answer by rothauserc(4718)

3. x = z/2
****************************************
the second equation becomes
z/2 + y = z
y = z/2
***************************************
the first equation becomes
z/2 + z/2 + z = 90
2z = 90
z = 45
x = 22.5 and y = 22.5
***************************************
equation 1
22.5 + 22.5 + 45 = 90
90 = 90
************************************
equation 2
22.5 + 22.5 = 45
45 = 45
*************************************
equation 3
22.5 = 45/2
22.5 = 22.5
*************************************